### v1

# The infinitude of square-free palindromes

Abstract.
We settle an open problem regarding palindromes; that is, positive integers which are the same when written forwards and backwards. In particular, we prove that for any fixed base $b\geq 2$, there exist infinitely many square-free palindromes in base $b$. We also provide an asymptotic expression for the number of such integers $\leq x$. The core of our proof utilises a hybrid $p$-adic/Archimedean van der Corput process, used in conjunction with an equidistribution estimate of Tuxanidy and Panario, as well as an elementary argument of Cilleruelo, Luca and Shparlinski.
The first author was supported by Australian Research Council Discovery Project DP240100186 and an Australian Mathematical Society Lift-off Fellowship.
The second author was supported by Australian Research Council Grants DE220100859 and DP230100534
Dedicated to Igor Shparlinski on the occasion of his 70th birthday
footnotetext: Corresponding author: Daniel Johnston (daniel.johnston@unsw.edu.au).footnotetext: Affiliation: School of Science, The University of New South Wales Canberra, Australia.footnotetext: Key phrases: Palindromes, square-free numbers, van der Corput method, exponential sums.footnotetext: 2020 Mathematics Subject Classification: 11A63, 11L07 (Primary) 11N25 (Secondary).
1  Introduction
1.1  Overview
For a fixed $b\geq 2$, palindromic numbers, or simply palindromes, are positive integers that are the same when written forwards and backwards in base $b$. In base 10, examples include $494$ and $12321$.
Prior to the 21st century, studies of palindromic numbers were primarily confined to the realm of recreational mathematics. However, in recent years there has been an influx of deep analytic and number-theoretic results regarding palindromes. Examples include the work of Cilleruelo, Luca and Baxter on sums of palindromes [12], or the work of Tuxanidy and Panario [32] on almost-prime palindromes. See also  [25, 2, 3, 24, 15, 13, 14, 33] for other modern works on palindromes. Notably, the recent interest in palindromes is in tandem with other studies of the digital properties of numbers. This includes work on primes with missing digits [27, 28, 23], with preassigned digits [7, 8, 30] and reversed primes [4, 16, 5, 17].
Despite the recent interest in palindromes, little is known about their multiplicative structure. One of the driving reasons for this is that palindromes are so sparse: for a fixed base $b$, there are only about $\sqrt{x}$ palindromes less than $x$. For reference, it is not known whether the similarly sized set
$S(x)=[x,x+\sqrt{x}],$
always contains a prime for sufficiently large $x$, even under assumption of the Riemann hypothesis (see e.g. [9, §1.2]).
The problem of whether there exists infinitely many square-free palindromes has often been mentioned in the literature (e.g. [3, p. 10], [24, p. 487], [16, p. 7], [11, p. 371]) and appears as part of Problem 58 in Shparlinski’s list of Problems on Exponential and Character Sums [29]. Here, as usual, a square-free number is a positive integer with no repeated prime factors. In this paper we resolve this problem, and in fact give an asymptotic for the number of square-free palindromes satisfying a standard coprimality condition (see Theorem 1.2).
Theorem 1.1.
For all $b\geq 2$, there exists infinitely many square-free palindromes in base $b$.
In the context of Theorem 1.1, it is worth noting that Carillo–Santana [10, Theorem 1.3], as well as Chourasiya and the first author [11, Theorem 1.6], recently and independently proved the existence (and an asymptotic) for cube-free palindromes. However, in both works, it is clear that their techniques are not strong enough to detect square-free palindromes. In another vein, Tuxanidy and Panario recently showed that there are infinitely many palindromes with at most $6$ prime factors [32, Theorem 1.4]. Although this latter result is very deep, their methods are again not strong enough to detect square-free palindromes. In particular, they do not provide a way to bound the number of palindromes with a large square divisor.
To prove Theorem 1.1, we thereby introduce a new approach which combines an elementary argument of Cilleruelo, Luca and Shparlinski [13] with the van der Corput method of exponential sums. Throughout our proof of Theorem 1.1, we prove a series of exponential sum and Fourier-analytic estimates, and new bounds for the number of palindromes with a large square divisor. These subsidiary results may be of independent interest, particularly in the use of weighted sieve methods applied to sets containing palindromes (see [32, § 11]).
Finally we note that Theorem 1.1 can be viewed as an approximation of the conjecture that there exists infinitely many palindromes which are also prime numbers [^1] In private communication, C. Maistret suggested using the term primendromes to succinctly refer to such numbers.. A proof of this conjecture still remains firmly out of reach, however it is known that almost all palindromes are composite [2].
1.2  Main notation and the asymptotic result
Let $b\geq 2$ be a fixed base. To more precisely define base-$b$ palindromes, consider the base $b$ expansion of an $N$-digit integer $n$
$n=\sum_{0\leq i<N}n_{i}b^{i},\quad 0\leq n_{i}<b.$ (1.1)
If $n_{N-1}\neq 0$ and $n_{0}\neq 0$ we then define the digital reverse of $n$, denoted $\overleftarrow{n}$, as
$\overleftarrow{n}=\sum_{0\leq i<N}n_{i}b^{N-1-i}.$ (1.2)
The set of base-$b$ palindromes is then succinctly defined as
$\mathscr{P}_{b}=\{n\geq 0:b\nmid n\ \text{and}\ n=\overleftarrow{n}\}$
with the condition $b\nmid n$ encapsulating the need for $n_{0}\neq 0$ in our expressions (1.1) and (1.2). We also define
$\Pi_{b}(N):=\mathscr{P}_{b}\cap[b^{N-1},b^{N})\quad\text{and}\quad\mathscr{P}_{b}(x):=\{n\in\mathscr{P}_{b}:n\leq x\}$
for the set of $N$-digit palindromes and the set of palindromes $\leq x$ respectively.
When studying palindromes, it is common to restrict to palindromes $n$ with $(n,b^{3}-b)=1$. This is to avoid technicalities owing to relations between $n$ and $\overleftarrow{n}$ mod $b^{3}-b$; see the introduction of [4] for further discussion on this matter. Consequently, we will work with the restricted sets
$\Pi_{b}^{*}(N)$ $:=\{n\in\Pi_{b}(N):(n,b^{3}-b)=1\},$
$\mathscr{P}_{b}^{*}(x)$ $:=\{n\in\mathscr{P}_{b}(x):(n,b^{3}-b)=1\}.$ (1.3)
Since an $N$-digit palindrome is completely determined by its first $\lceil N/2\rceil$ digits, a simple combinatorial argument yields that
$\Pi_{b}(N)\asymp_{b}b^{N/2}\quad\text{and}\quad\#\mathscr{P}_{b}(x)\asymp_{b}\sqrt{x}.$
Similarly, one can show (see [32, Lemma 9.1])
$\Pi_{b}^{*}(N)\asymp_{b}b^{N/2}\quad\text{and}\quad\#\mathscr{P}_{b}^{*}(x)\asymp_{b}\sqrt{x}.$ (1.4)
With the above notation, an asymptotic version of our result (Theorem 1.1) is given as follows.
Theorem 1.2.
For all bases $b\geq 2$, if
$Q_{b}^{*}(x):=\#\{n\in\mathscr{P}_{b}^{*}(x):n\ \text{is square-free}\}$
then one has, for any $A>0$,
$Q_{b}^{*}(x)=\frac{\#\mathscr{P}_{b}^{*}(x)}{\zeta(2)}\prod_{p\mid b^{3}-b}\left(1-\frac{1}{p^{2}}\right)^{-1}\left(1+O_{b,A}\left(\frac{1}{(\log x)^{A}}\right)\right),$ (1.5)
where $\zeta(\cdot)$ is the Riemann zeta function.
Notably, the expression in (1.5) has the same form as the asymptotic expression for $k$-free palindromes ($k\geq 3$) given in [11, Theorem 1.6]. Here,
$\frac{1}{\zeta(2)}\prod_{p\mid b^{3}-b}\left(1-\frac{1}{p^{2}}\right)^{-1}$
is the density of square-free numbers that are coprime to $b^{3}-b$. Therefore, (1.5) essentially says that after restricting to integers $n$ with $(n,b^{3}-b)=1$, the probability of $n$ being a palindrome is independent of the probability of $n$ being square-free.
The behaviour of square-free palindromes in the unrestricted set $\mathscr{P}_{b}(x)$ appears to be much more chaotic. However, if we fix the digital length $N$ and consider
$\mathscr{Q}_{b}(N):=\#\{n\in\Pi_{b}(N):n\ \text{is square-free}\}$
for the number of square-free elements of $\Pi_{b}(N)$ instead of $\mathscr{P}_{b}^{*}(x)$, then numerical computations suggest that
$\mathscr{Q}_{b}(N)\sim\frac{\#\Pi_{b}(N)}{\zeta(2)}.$ (1.6)
Whilst the expression (1.6) is very elegant, we do not attempt to prove it here. This is because working with $\mathscr{P}_{b}^{*}(x)$ or $\Pi_{b}^{*}(N)$ rather than $\mathscr{P}_{b}(x)$ or $\Pi_{b}(N)$ allows us to simplify many arguments. However, we believe that it should be possible to prove (1.6) by a routine reworking of our approach and the results we rely on from the literature.
1.3  Structure of paper
A summary of the rest of the paper is as follows. In Section 2 we give an outline of the proof of Theorem 1.2 and thus Theorem 1.1. More precisely, we state our main subsidiary results (Propositions 2.2, 2.3 and 2.4) and show how these imply Theorem 1.2. In Section 3 we then prove the first of these subsidiary results, Proposition 2.2, which gives an elementary bound on the number of palindromes with a large square divisor. Next, in Section 4 we discuss and prove all of the preliminaries required for our variant of the van der Corput method. Finally, we conclude in Sections 5 and 6 by giving the proofs of Propositions 2.3 and 2.4.
Throughout we employ Vinogradov’s notation $f\ll g$ to mean $f=O(g)$. We also use the standard notation $\mu(d)$ for the Möbius function, $e(x):=\exp(2\pi ix)$ for complex exponentials, and $d\sim D$ to mean that $d\in[D,2D]$. This latter notation should not be confused with the identical notation $f(x)\sim g(x)$ (see e.g. (1.6)) reserved for an asymptotic equivalence of functions.
2  Outline of the proof of Theorem 1.2
To set up the proof of Theorem 1.2, we first use Möbius inversion to write
$Q_{b}^{*}(x)=\sum_{n\in\mathscr{P}_{b}^{*}(x)}\mu^{2}(n)=\sum_{n\in\mathscr{P}_{b}^{*}(x)}\sum_{d^{2}\mid n}\mu(d)=\sum_{\begin{subarray}{c}d\leq\sqrt{x}\\ (d,b^{3}-b)=1\end{subarray}}\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(x)\\ d^{2}\mid n\end{subarray}}\mu(d).$ (2.1)
Our underlying approach is then to split the sum in (2.1) into different ranges for $d$ and apply relevant estimates. A formal proof of how these estimates imply Theorem 1.2 is given at the end of this section. The first estimate we use is the following equidistribution result for square moduli, proven by Tuxanidy and Panario [32] as an application of a large sieve inequality of Baier and Zhao [1].
Proposition 2.1 ([32, Proposition 10.1]).
For any $b\geq 2$, $\varepsilon>0$ and $A>0$, one has
$\sum_{\begin{subarray}{c}d\leq x^{1/4-\varepsilon}\\ (d,b^{3}-b)=1\end{subarray}}\mu^{2}(d)\sup_{y\leq x}\max_{a\in\mathbb{Z}}\left|\sum_{n\in\mathscr{P}_{b}^{*}(y)}\left(\mathbb{1}_{n\equiv a\mathrm{(}\mathrm{mod}\ d^{2}\mathrm{)}}-\frac{1}{d^{2}}\right)\right|\ll_{A,b,\varepsilon}\frac{\#\mathscr{P}_{b}^{*}(x)}{(\log x)^{A}}.$ (2.2)
Proposition 2.1 allows us to treat the range $1\leq d\leq x^{1/4}$ in (2.1), from which we can extract the main term of $Q_{b}^{*}(x)$. After this, we need sufficiently strong upper bounds for the number of palindromes which have a large square divisor $d^{2}$, with $d>x^{1/4-\varepsilon}$. This will be done by considering the function
$S_{b}(x,D):=\#\{n\in\mathscr{P}^{*}_{b}(x):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$
which counts the number of $n\in\mathscr{P}_{b}^{*}(x)$ divisible by $d^{2}$ with $d$ in a dyadic interval $[D,2D]$. The first bound we obtain for $S_{b}(x,D)$ is given below. Recalling that $\mathscr{P}_{b}^{*}(x)\asymp_{b}\sqrt{x}$ (see (1.4)), we find that this result is nontrivial in the range $x^{1/3+\varepsilon}\leq D\leq\sqrt{x}$.
Proposition 2.2.
For all $b\geq 2$ and $D\geq 1$, we have
$S_{b}(x,D)\ll_{b}\frac{x}{D^{3/2}}.$
In particular, for any $\varepsilon>0$, if $D\geq x^{1/3+\varepsilon}$ then
$S_{b}(x,D)\ll_{b}x^{1/2-\varepsilon}.$
The proof of Proposition 2.2 is relatively simple, being an elaboration of an elementary argument due to Cilleruelo, Luca and Shparlinski [13, Theorem 1] which reduces counting palindromes with a large square divisor to counting integers in short intervals and arithmetic progressions.
Given Proposition 2.2, it remains to cover the range $x^{1/4-\varepsilon}<D<x^{1/3+\varepsilon}$. To do so, we incorporate Fourier-analytic and exponential sum methods into the proof of Proposition 2.2, yielding bounds for $S_{b}(x,D)$ that are better for smaller values of $D$. The overarching approach here will be to apply a van der Corput method of exponential sums. For background on this method see [19]. An explanation of our modification of this method is then given in Section 4. Ultimately, we are able to obtain the following two bounds, which are non-trivial for different ranges of $D$.
Proposition 2.3.
For all $b\geq 2$, $x^{1/4}\leq D\leq x^{2/5}$ and $\varepsilon>0$, we have
$S_{b}(x,D)\ll_{b,\varepsilon}\frac{x^{2/3}}{D^{2/3-\varepsilon}}.$ (2.3)
In particular, for any $\varepsilon\in(0,0.25)$, if $x^{1/4+3\varepsilon}\leq D\leq x^{2/5}$ then
$S_{b}(x,D)\ll_{b,\varepsilon}x^{1/2-\varepsilon}.$
Proposition 2.4.
For all $b\geq 2$ and $\varepsilon>0$, if $x^{3/13}\leq D\leq x^{8/31-\varepsilon}$ then
$S_{b}(x,D)\ll_{b,\varepsilon}\frac{x^{7/11}}{D^{13/22-\varepsilon}}.$ (2.4)
In particular, for any $\varepsilon\in(0,0.1)$, if $x^{3/13+3\varepsilon}\leq D\leq x^{8/31-\varepsilon}$ then
$S_{b}(x,D)\ll_{b,\varepsilon}x^{1/2-\varepsilon}.$
Combining Propositions 2.2, 2.3 and 2.4 yields the following theorem.
Theorem 2.5.
For all $b\geq 2$ and $\varepsilon>0$, if $D\geq x^{3/13+\varepsilon}$ then there exists a $\delta>0$ such that
$S_{b}(x,D)\ll_{b,\varepsilon}x^{1/2-\delta}.$
Since $3/13<1/4$, Theorem 2.5 is sufficient for our purposes.
We conclude this section by formalising the above discussion, giving an explicit proof as for how Theorem 1.2 follows from Proposition 2.1 and Theorem 2.5. The rest of the paper is then concerned with the proofs of Propositions 2.2, 2.3 and 2.4.
Proof of Theorem 1.2 assuming Proposition 2.1 and Theorem 2.5..
Beginning with (2.1), we write
$Q_{b}^{*}(x)=\sum_{\begin{subarray}{c}d\leq\sqrt{x}\\ (d,b^{3}-b)=1\end{subarray}}\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(x)\\ d^{2}\mid n\end{subarray}}\mu(d)=S_{1}(x)+S_{2}(x),$
where
$S_{1}(x)$ $:=\sum_{\begin{subarray}{c}d\leq x^{0.24}\\ (d,b^{3}-b)=1\end{subarray}}\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(x)\\ d^{2}\mid n\end{subarray}}\mu(d),$
$S_{2}(x)$ $:=\sum_{\begin{subarray}{c}x^{0.24}<d\leq\sqrt{x}\\ (d,b^{3}-b)=1\end{subarray}}\>\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(x)\\ d^{2}\mid n\end{subarray}}\mu(d).$
We first deal with $S_{1}(x)$. Here, Proposition 2.1 implies that
$S_{1}(x)$ $=\#\mathscr{P}_{b}^{*}(x)\sum_{\begin{subarray}{c}d\leq x^{0.24}\\ (d,b^{3}-b)=1\end{subarray}}\frac{\mu(d)}{d^{2}}+O_{A,b}\left(\frac{\#\mathscr{P}_{b}^{*}(x)}{(\log x)^{A}}\right)$
$=\#\mathscr{P}_{b}^{*}(x)\left(\sum_{\begin{subarray}{c}d\geq 1\\ (d,b^{3}-b)=1\end{subarray}}\frac{\mu(d)}{d^{2}}-\sum_{\begin{subarray}{c}d>x^{0.24}\\ (d,b^{3}-b)=1\end{subarray}}\frac{\mu(d)}{d^{2}}\right)+O_{A,b}\left(\frac{\#\mathscr{P}_{b}^{*}(x)}{(\log x)^{A}}\right).$ (2.5)
Converting the first sum in (2.5) into an Euler product gives
$\sum_{\begin{subarray}{c}d\geq 1\\ (d,b^{3}-b)=1\end{subarray}}\frac{\mu(d)}{d^{2}}=\prod_{\begin{subarray}{c}p\geq 2\\ (p,b^{3}-b)=1\end{subarray}}\left(1-\frac{1}{p^{2}}\right)=\frac{1}{\zeta(2)}\prod_{p\mid b^{3}-b}\left(1-\frac{1}{p^{2}}\right)^{-1}.$ (2.6)
Then,
$\sum_{\begin{subarray}{c}d>x^{0.24}\\ (d,b^{3}-b)=1\end{subarray}}\frac{\mu(d)}{d^{2}}\ll\sum_{d>x^{0.24}}\frac{1}{d^{2}}\ll\frac{1}{x^{0.24}}.$ (2.7)
Hence, upon substituting (2.6) and (2.7) into (2.5), one has
$S_{1}(x)=\frac{\#\mathscr{P}_{b}^{*}(x)}{\zeta(2)}\prod_{p\mid b^{3}-b}\left(1-\frac{1}{p^{2}}\right)^{-1}\left(1+O_{A,b}\left(\frac{1}{(\log x)^{A}}\right)\right).$
This is the desired asymptotic for $Q_{b}^{*}(x)$ in Theorem 1.2. Therefore, it suffices to show that $S_{2}(x)$ is of a smaller order. With this aim, we apply the triangle inequality and split into dyadic intervals to yield
$|S_{2}(x)|\leq\sum_{x^{0.24}\leq d\leq\sqrt{x}}\>\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(x)\\ d^{2}\mid n\end{subarray}}1$ $\leq\sum_{m=\lfloor\log_{2}x^{0.24}\rfloor}^{\lceil\log_{2}\sqrt{x}\rceil}\>\sum_{d\sim 2^{m}}\>\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(x)\\ d^{2}\mid n\end{subarray}}1$
$=\sum_{m=\lfloor\log_{2}x^{0.24}\rfloor}^{\lceil\log_{2}\sqrt{x}\rceil}S_{b}(x,2^{m}),$
where we have used the standard notation $\log_{2}x:=\log x/\log 2$. Applying Theorem 2.5 then gives some $\delta>0$ such that
$S_{2}(x)\ll_{b}\sum_{m=\lfloor\log_{2}x^{0.24}\rfloor}^{\lceil\log_{2}\sqrt{x}\rceil}x^{\frac{1}{2}-\delta}\ll x^{\frac{1}{2}-\delta}\log x=o_{b,A}\left(\frac{\#\mathscr{P}_{b}^{*}(x)}{(\log x)^{A}}\right)$
as desired. ∎
3  Proof of Proposition 2.2
In this section we prove Proposition 2.2. The proof is elementary and inspired by that of [13, Theorem 1]. In particular, the simplest case of [13, Theorem 1] gives an upper bound for the number of square palindromes and we adapt their proof to obtain an upper bound for the number of palindromes with a large square divisor.
To begin with, we give a couple of simple lemmas.
Lemma 3.1.
Suppose for some $d,X,Y\geq 1$ that
$|d^{2}-Y|\leq X.$
Then,
$|d-\sqrt{Y}|\leq\frac{X}{\sqrt{Y}}.$
Proof.
Since $d,Y\geq 1$, we have
$\sqrt{Y}|d-\sqrt{Y}|\leq|(d+\sqrt{Y})(d-\sqrt{Y})|=|d^{2}-Y|\leq X$
from which the result follows. ∎
Lemma 3.2.
Let $a\in\mathbb{Z}$ and $q\geq 2$. Then, the number of solutions $w$ mod $q$ with $(w,q)=1$ to the congruence
$w^{k}\equiv a\pmod{q}$ (3.1)
is less than $k^{\omega(q)}$, where $\omega(q)$ is the number of distinct prime factors of $q$.
Proof.
The case where $q$ is a prime power is, for example, covered in Section 4.2 of the textbook [22], where one sees that (3.1) has at most $k$ solutions. For general $q$, one then applies the Chinese remainder theorem. ∎
We now prove Proposition 2.2.
Proof of Proposition 2.2.
We begin by proving the result for palindromes of fixed digit length $N$. That is, we show
$\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}\ll_{b}\frac{b^{N}}{D^{3/2}}.$
Let $K$ be the largest integer such that $b^{K}\leq\sqrt{D}$. This implies that
$b^{K}\asymp_{b}\sqrt{D}.$ (3.2)
Now, there exists an integer $a$ with $0\leq a<b^{K}$ such that
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$
$\qquad\qquad\leq b^{K}\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n,\ n\equiv a\ \text{(mod}\ b^{K}\text{)}\}.$ (3.3)
Let $n\in\Pi_{b}^{*}(N)$ be a palindrome counted in (3). The condition $n\equiv a\pmod{b^{K}}$ indicates that the least significant $K$ digits of $n$ are equal to $a$. Since the last base-$b$ digit of a palindrome cannot be 0, we may assume that $a$ does not end in a 0. Now, let $\alpha$ be the $N$-digit number with its leading digits the mirror image of $a$ followed by trailing zeros. So in particular, $\alpha$ determines the $K$ most significant digits of the palindrome $n$. We thus have
$\left|n-\alpha\right|\leq\frac{b^{N}}{b^{K}}.$
Therefore,
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$
$\leq b^{K}\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n,\ n\equiv a\ \text{(mod}\ b^{K}\text{)}\}$
$\leq b^{K}\#\left\{d\sim D,\ s\in\left[\frac{b^{N-1}}{4D^{2}},\frac{b^{N}}{D^{2}}\right]:d^{2}s\equiv a\ \text{(mod}\ b^{K}\text{)},\ (sd,b)=1,\ \left|d^{2}s-\alpha\right|\leq\frac{b^{N}}{b^{K}}\right\}$
where we have the condition $(sd,b)=1$ since any $n\in\Pi_{b}^{*}(N)$ is coprime to $b$. From here, we deduce that there exists $s\in[b^{N-1}/4D^{2},b^{N}/D^{2}]$ with $(s,b)=1$ and
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$
$\quad\ll\frac{b^{N+K}}{D^{2}}\#\left\{d\sim D:d^{2}\equiv as^{-1}\ \text{(mod}\ b^{K}\text{)},\ (d,b)=1,\ \left|d^{2}s-\alpha\right|\leq\frac{b^{N}}{b^{K}}\right\}.$ (3.4)
By Lemma 3.1, we have
$\left|d^{2}s-\alpha\right|\leq\frac{b^{N}}{b^{K}}\Rightarrow\left|d-\sqrt{\frac{\alpha}{s}}\right|\ll_{b}\frac{D}{b^{K}}$ (3.5)
where we have used $\alpha\asymp_{b}b^{N}$ and $s\asymp_{b}b^{N}/D^{2}$. Now, since $(d,b)=1$, Lemma 3.2 implies that the number of $d$ in an interval $[X,Y]$ with $d^{2}\equiv as^{-1}$ (mod $b^{K}$) is $\ll(Y-X)/b^{K}+1$. Hence, the number of $d$ satisfying (3.5) and $d^{2}\equiv as^{-1}\ \text{(mod $b^{K}$)}$} is $\ll D/b^{2K}+1$. Substituting this information into (3) and applying (3.2) yields
$\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}\ll_{b}\frac{b^{N}}{Db^{K}}+\frac{b^{N+K}}{D^{2}}\ll_{b}\frac{b^{N}}{D^{3/2}}$
as desired. To convert this to a result over all $n\in\mathscr{P}_{b}^{*}(x)$ we simply note that
$\sum_{d\sim D}\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(x)\\ d^{2}\mid n\end{subarray}}1=\sum_{d\sim D}\sum_{N=1}^{\lceil\log x/\log b\rceil}\sum_{\begin{subarray}{c}n\in\Pi_{b}^{*}(N)\\ d^{2}\mid n\end{subarray}}1\ll_{b}\sum_{N=1}^{\lceil\log x/\log b\rceil}\frac{b^{N}}{D^{3/2}}\ll_{b}\frac{x}{D^{3/2}},$
as required. ∎
4  Preliminaries for the van der Corput method
To prove Propositions 2.3 and 2.4, we modify the proof of Proposition 2.2 by applying a hybrid $p$-adic and Archimedean van der Corput method to detect integers in arithmetic progressions and short intervals. This method uses two processes:
  • An $A$-process which is based on Weyl differencing (see Lemma 4.1).
  • A $B$-process which is based on Poisson summation (see Lemma 4.2).
The proof of Proposition 2.3 uses only the $B$-process which sets the “convexity barrier” $D\geq x^{1/4+\varepsilon}$}. Then, Proposition 2.4 uses a combination of the $B$-process and the $A$-process to treat slightly smaller values of $D$ and break the $x^{1/4}$-barrier. One could certainly try to elaborate on our approach to obtain non-trivial results for even smaller $D$, however this is unnecessary for our application.
Throughout the rest of this section, we provide all of the preliminaries required for this van der Corput method. Namely, we state the $A$ and $B$ processes we use, and prove a series of relevant results on oscillatory integrals and exponential sums.
4.1  The $A$ and $B$ processes
The $A$-process we use is a smooth variation of the usual Weyl-van der Corput inequality. The proof is similar to that of the unsmoothed inequality (see e.g. [19, Lemma 2.5]). However, we include all the details here for clarity.
Lemma 4.1.
Let $z_{n}$ be a sequence of complex numbers satisfying $|z_{n}|=O(1)$, and $\psi$ a smooth function with support in $[1/2,5/2]$ and $\psi(x)=1$ for $x\in[1,2]$}. Then for any $Q\leq D^{1/2}$,
$\left|\sum_{d\sim D}z_{d}\right|\ll\frac{D}{Q^{1/2}}+\frac{D^{1/2}}{Q^{1/2}}\left(\sum_{q\leq Q}\left|\sum_{d\in\operatorname{\mathbb{Z}}}\psi\left(\frac{d}{D}\right)z_{d}\overline{z}_{d+q}\right|\right)^{1/2}.$
Proof.
Since for any $q\leq Q$
$\sum_{d\sim D}z_{d}=\sum_{d+q\sim D}z_{d+q}=\sum_{d\sim D}z_{d+q}+O(Q)$
we have
$\sum_{d\sim D}z_{d}$ $=\frac{1}{Q}\sum_{\begin{subarray}{c}1\leq q\leq Q\end{subarray}}\sum_{\begin{subarray}{c}d\sim D\end{subarray}}z_{d+q}+O(Q)\ll\frac{1}{Q}\sum_{d\sim D}\left|\sum_{\begin{subarray}{c}1\leq q\leq Q\end{subarray}}z_{d+q}\right|+Q.$
Hence, by the Cauchy-Schwarz inequality,
$\left|\sum_{d\sim D}z_{d}\right|^{2}\ll\frac{D}{Q^{2}}\sum_{d\sim D}\left|\sum_{\begin{subarray}{c}1\leq q\leq Q\end{subarray}}z_{d+q}\right|^{2}\ll\frac{D}{Q^{2}}\sum_{d\in\operatorname{\mathbb{Z}}}\psi\left(\frac{d}{D}\right)\left|\sum_{\begin{subarray}{c}1\leq q\leq Q\end{subarray}}z_{d+q}\right|^{2}+Q^{2}$ (4.1)
with $\psi$ as in the statement of the lemma. Expanding the square in (4.1) gives
$\left|\sum_{d\sim D}z_{d}\right|^{2}\ll\frac{D^{2}}{Q}+\frac{D}{Q^{2}}\sum_{\begin{subarray}{c}1\leq q_{1}<q_{2}\leq Q\end{subarray}}\left|\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\end{subarray}}\psi\left(\frac{d}{D}\right)z_{d+q_{1}}\overline{z}_{d+q_{2}}\right|+Q^{2},$ (4.2)
with the $D^{2}/Q$ term coming from removing the diagonal contribution with $q_{1}=q_{2}$. Next, we perform a change of variable $d\to d-q_{1}$ so that
$\left|\sum_{d\sim D}z_{d}\right|^{2}\ll\frac{D^{2}}{Q}+Q^{2}+\frac{D}{Q^{2}}\sum_{\begin{subarray}{c}1\leq q_{1}<q_{2}\leq Q\end{subarray}}\left|\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\end{subarray}}\psi\left(\frac{d-q_{1}}{D}\right)z_{d}\overline{z}_{d+(q_{2}-q_{1})}\right|.$
By the mean value theorem
$\psi\left((d-q_{1}/D\right)=\psi(d/D)+O(Q/D).$
Hence, since $Q\leq D^{1/2}$,
$\left|\sum_{d\sim D}z_{d}\right|^{2}$ $\ll\frac{D^{2}}{Q}+DQ+\frac{D}{Q^{2}}\sum_{1\leq q_{1}<q_{2}\leq Q}\left|\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\end{subarray}}\psi\left(\frac{d}{D}\right)z_{d}\overline{z}_{d+(q_{2}-q_{1})}\right|$
$\ll\frac{D^{2}}{Q}+\frac{D}{Q^{2}}\sum_{1\leq q_{1}<q_{2}\leq Q}\left|\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\end{subarray}}\psi\left(\frac{d}{D}\right)z_{d}\overline{z}_{d+(q_{2}-q_{1})}\right|,$ (4.3)
where in the second line we have used that $DQ\leq D^{2}/Q$ since $Q\leq D^{1/2}$. Finally, we write $q=q_{2}-q_{1}$ and note that each difference $q_{2}-q_{1}$ occurs at most $Q$ times in (4.3), and the desired result follows. ∎
The $B$-process is then given by the following variation of Poisson summation which, for example, may be readily derived from [6, Lemma 2.9].
Lemma 4.2.
Let $f\in L^{1}(\operatorname{\mathbb{R}})$ be a continuous, compactly supported function, and $g$ be a periodic function mod $q$. Then
$\sum_{n\in\operatorname{\mathbb{Z}}}f(n)g(n)=\frac{1}{q^{1/2}}\sum_{m\in\operatorname{\mathbb{Z}}}\widehat{f}\left(\frac{m}{q}\right)\widehat{g}(m)$ (4.4)
[…]
$\widehat{f}(k)=\int_{\mathbb{R}}f(u)e(-ku)\mathrm{d}u$ (4.5)
and $\widehat{g}$ denotes the Fourier transform of $g$ in $\operatorname{\mathbb{Z}}/q\operatorname{\mathbb{Z}}$:
$\widehat{g}(m)=\frac{1}{q^{1/2}}\sum_{y=1}^{q}g(y)e\left(\frac{my}{q}\right).$
[…]
4.2  Oscillatory integral estimates
[…]
Lemma 4.3 ([20, Lemma 4]).
Let $W$ and $R$ be smooth functions on $[a,b]$. Let
$\Phi=\inf\{|\phi^{\prime}(x)|\ :x\in[a,b]\}.$
Suppose further that
$W^{(k)}(x)\ll(b-a)^{k}$
and
$\phi^{(k)}(x)\ll(b-a)^{1-k}\Phi.$
Then for any $N\geq 1$
$\int_{a}^{b}e^{i\phi(x)}W(x)\mathrm{d}x\ll(b-a)^{1-N}\Phi^{-N}.$
We also require some results regarding the mean value of definite integrals. We begin with a statement of the “second mean value theorem”. See [21] for further discussion and a proof of this classical result.
Lemma 4.4 (Second mean value theorem).
Let $g:[a,b]\to\operatorname{\mathbb{R}}$ be integrable. If $f:[a,b]\to\operatorname{\mathbb{R}}$ is a positive monotonically decreasing function, then there exists $c\in(a,b]$ such that
$\int_{a}^{b}f(x)g(x)\mathrm{d}x=f(a^{+})\int_{a}^{c}g(x)\mathrm{d}x,$
where $f(a^{+})=\lim_{x\to a^{+}}f(x)$. On the other hand, if $f:[a,b]\to\operatorname{\mathbb{R}}$ is a positive monotonically increasing function, then there exists $c\in[a,b)$ such that
$\int_{a}^{b}f(x)g(x)\mathrm{d}x=f(b^{-})\int_{c}^{b}g(x)\mathrm{d}x,$
where $f(b^{-})=\lim_{x\to b^{-}}f(x)$.
Using the second mean value theorem, we state two well known estimates on oscillatory integrals. Similar results are given in the literature (e.g. [31, Lemmas 4.3 and 4.5]); however they are not in the right form for our purposes so we provide complete details.
Lemma 4.5.
Let $F(x)$ be a differentiable monotonic function on $[a,b]$ such that $F^{\prime}(x)>m>0$} or $F^{\prime}(x)<-m<0$. Suppose also that $G(x)$ is a montonic function on the interval $[a,b]$ with $0\leq G(x)\leq M$ for some constant $M>0$. Then,
$\left|\int_{a}^{b}G(x)e^{iF(x)}\mathrm{d}x\right|\leq\frac{4M}{m}.$ (4.6)
Proof.
First we consider the real part of the integral
$\int_{a}^{b}G(x)\cos(F(x))\mathrm{d}x.$ (4.7)
Throughout we assume that $F^{\prime}(x)$ and $G(x)$ are monotonically increasing on $[a,b]$, since the argument is essentially identical for the other cases of monotonicity. By the second mean value theorem (Lemma 4.4), there exists $c\in[a,b)$ such that
$\int_{a}^{b}G(x)\cos(F(x))\mathrm{d}x=G(b)\int_{c}^{b}\cos(F(x))\mathrm{d}x.$
If $F^{\prime}(x)>m>0$, then applying the second mean value theorem again, we have that there exists $d\in(c,b]$ such that
$G(b)\int_{c}^{b}\cos(F(x))\mathrm{d}x$ $=\frac{G(b)}{F^{\prime}(c)}\int_{c}^{d}F^{\prime}(x)\cos(F(x))\mathrm{d}x$
$=\frac{G(b)}{F^{\prime}(c)}\left(\sin\{F(d)\}-\sin\{F(c)\}\right).$
Otherwise, if $F^{\prime}(x)<-m<0$, then $1/F^{\prime}(x)$ is positive and increasing and there exists $e\in[c,b)$ such that
$G(b)\int_{c}^{b}\cos(F(x))\mathrm{d}x$ $=-\frac{G(b)}{F^{\prime}(b)}\int_{e}^{b}F^{\prime}(x)\cos(F(x))\mathrm{d}x$
$=-\frac{G(b)}{F^{\prime}(b)}\left(\sin\{F(b)\}-\sin\{F(e)\}\right).$
In either case,
$\left|\int_{a}^{b}G(x)\cos(F(x))\mathrm{d}x\right|=\left|G(b)\int_{c}^{b}\cos(F(x))\mathrm{d}x\right|\leq\frac{2M}{m}.$
The same argument holds for the imaginary part of the integral in (4.6), namely,
$\left|\int_{a}^{b}G(x)\sin(F(x))\mathrm{d}x\right|\leq\frac{2M}{m}.$
Hence, adding the bounds for real and imaginary parts together we obtain the desired result. ∎
Lemma 4.6.
Let $F(x)$ be a real, twice differentiable function with either $F^{\prime\prime}(x)>r>0$} or $F^{\prime\prime}(x)<-r<0$ throughout the interval $[a,b]$. Suppose also that $G(x)$ is a real piecewise montonic function on the interval $[a,b]$ with $0\leq G(x)\leq M$ for some constant $M>0$. Then,
$\left|\int_{a}^{b}G(x)e^{iF(x)}\mathrm{d}x\right|\leq\frac{8KM}{\sqrt{r}}\ll_{K,M}\frac{1}{\sqrt{r}},$ (4.8)
where $K$ is the number of piecewise monotonic components of $G$.
Proof.
We begin by making a finite subdivision of the integral in (4.8) as
$\int_{a}^{b}G(x)e^{iF(x)}\mathrm{d}x=\sum_{k=1}^{K}\int_{a_{k}}^{b_{k}}G(x)e^{iF(x)}\mathrm{d}x$ (4.9)
so that $G(x)$ is monotonic on each $a_{k},b_{k}]$ except possibly at the endpoints.
Fix $k\in\{1,\ldots,K\}$. From here onwards we assume $G(x)$ is monotonically increasing on $a_{k},b_{k}]$ and $F^{\prime\prime}(x)>r>0$ (so that $F^{\prime}(x)$ is increasing) since the proofs of the other cases are essentially the same. So, since $G(x)$ is monotonically increasing on $a_{k},b_{k}]$, we have by the second mean value theorem (Lemma 4.4) that there exists $c_{k}\in[a_{k},b_{k})$ such that
$\left|\int_{a_{k}}^{b_{k}}G(x)e^{iF(x)}\mathrm{d}x\right|\leq M\left|\int_{c_{k}}^{b_{k}}e^{iF(x)}\mathrm{d}x\right|.$ (4.10)
Then, since $F^{\prime}(x)$ is increasing on $c_{k},b_{k}]$ it has at most one zero in this interval, which we denote by $d_{k}$. If $F^{\prime}(x)$ does not vanish on $c_{k},b_{k}]$ then we just let $d_{k}$ be the midpoint $d_{k}=(b_{k}-c_{k})/2$. We now write
$\int_{c_{k}}^{b_{k}}e^{iF(x)}\mathrm{d}x=\int_{c_{k}}^{d_{k}-\delta}+\int_{d_{k}-\delta}^{d_{k}+\delta}+\int_{d_{k}+\delta}^{b_{k}}=I_{1}+I_{2}+I_{3},$ (4.11)
where $\delta>0$ is a small parameter which we choose later, and we assume that $d_{k}-\delta>c_{k}$ and $d_{k}+\delta<b_{k}$ for otherwise we can just remove $I_{1}$ or $I_{3}$ from (4.11). First we note that
$|I_{2}|\leq 2\delta$
since $|e^{iF(x)}|=1$. Now, since $F^{\prime}(x)$ is increasing, we have for any $x\in[d_{k}+\delta,b_{k}]$
$F^{\prime}(x)\geq F(d_{k}+\delta)=\int_{d_{k}}^{d_{k}+\delta}F^{\prime\prime}(t)\mathrm{d}t\geq r\delta.$
Thus, Lemma 4.5 gives
$|I_{3}|\leq\frac{4}{r\delta}$
and by analogous reasoning one obtains the same bound for $|I_{1}|$. Substituting these bounds into (4.11) and (4.10) one has
$\left|\int_{a_{k}}^{b_{k}}G(x)e^{iF(x)}\mathrm{d}x\right|\leq\frac{8M}{r\delta}+2M\delta.$
The desired result then follows upon setting $\delta=2/\sqrt{r}$ and summing over all $k\in\{1,\ldots,K\}$}. ∎
4.3  Complete exponential sum estimates
When applying the van der Corput method, we will need to estimate complete exponential sums of the following form:
$K_{2}(a_{1},a_{2},a_{3},q,c):=\frac{1}{c^{1/2}}\sum_{\begin{subarray}{c}1\leq x\leq c\\ (x(x+q),c)=1\end{subarray}}e\left(\frac{a_{1}x+a_{2}\overline{x}^{2}+a_{3}\overline{(x+q)}^{2}}{c}\right)$ (4.12)
where each argument of $K_{2}$ is a positive integer and $\overline{x}$ denotes the multiplicative inverse of $x$ modulo $c$. When $a_{3}=0$ and $q=0$ we simplify notation:
$K_{2}(a_{1},a_{2},c):=\frac{1}{c^{1/2}}\sum_{\begin{subarray}{c}1\leq x\leq c\\ (x,c)=1\end{subarray}}e\left(\frac{a_{1}x+a_{2}\overline{x}^{2}}{c}\right).$ (4.13)
Such sums may be referred to as “quadratic Kloosterman sums”. To bound (4.12) and (4.13), or averages thereof, we begin with the following lemma which is based on the method of $p$-adic stationary phase. We remark that a similar approach was used in the work of Mangerel [26], wherein estimates for sums of the form (4.13) were used to study square-free numbers in arithmetic progressions to smooth moduli.
Lemma 4.7.
Let $c$ be an integer with prime factorization
$c=p_{1}^{\alpha_{1}}\dots p_{k}^{\alpha_{k}}.$
For each $1\leq j\leq k$ write
$\alpha_{j,1}+\alpha_{j,2}=\alpha_{j},$
where $\alpha_{j,1}=\lfloor{\alpha_{j}/2}\rfloor$ and $\alpha_{j,2}=\lceil{\alpha_{j}/2}\rceil$. Define
$c_{1}=\prod_{j=1}^{k}p_{j}^{\alpha_{j,1}},\quad c_{2}=\prod_{j=1}^{k}p_{j}^{\alpha_{j,2}}.$
Then
$K_{2}(a_{1},a_{2},a_{3},q,c)$
$\quad\ll\#\{w\!\!\!\mod{c_{1}}:(w(w+q),c)=1,\ a_{1}-2a_{2}\overline{w}^{3}-2a_{3}\overline{(w+q)}^{3}\equiv 0\!\!\mod{c_{1}}\}$ (4.14)
with the implied constant depending on the primes $p_{1},\ldots,p_{k}$ but not $\alpha_{1},\ldots,\alpha_{k}$.
Proof.
We first note that any $1\leq x\leq c$ can be uniquely written as $x=w+zc_{2}$ with $1\leq w\leq c_{2}$ and $0\leq z<c_{1}$. Hence,
$K_{2}(a_{1},a_{2},a_{3},q,c)=\frac{1}{c^{1/2}}\sum_{\begin{subarray}{c}z\ \text{mod}\ {c_{1}}\\ w\ \text{mod}\ {c_{2}}\\ (w(w+q),c)=1\end{subarray}}e\left(\frac{F(w+zc_{2})}{c}\right)$
where
$F(x)=a_{1}x+a_{2}x^{\phi(c)-2}+a_{3}(x+q)^{\phi(c)-2},$
with $\phi$ the Euler totient function. Now, since $2\alpha_{j,2}\geq\alpha_{j}$ for each $1\leq j\leq k$, we have for any $m\geq 0$,
$(w+zc_{2})^{m}\equiv w^{m}+zc_{2}mw^{m-1}\pmod{c}$.
In particular,
$F(w+zc_{2})\equiv F(w)+zc_{2}F^{\prime}(w)\pmod{c}$
and thus
$K_{2}(a_{1},a_{2},a_{3},q,c)$ $=\frac{1}{c^{1/2}}\sum_{\begin{subarray}{c}z\ \text{mod}\ {c_{1}}\\ w\ \text{mod}\ {c_{2}}\\ (w(w+q),c)=1\end{subarray}}e\left(\frac{F(w)}{c}+\frac{zF^{\prime}(w)}{c_{1}}\right)$
$=\frac{1}{c^{1/2}}\sum_{\begin{subarray}{c}w\ \text{mod}\ {c_{2}}\\ (w(w+q),c)=1\end{subarray}}e\left(\frac{F(w)}{c}\right)\sum_{z\ \text{mod}\ {c_{1}}}\exp\left(\frac{zF^{\prime}(w)}{c_{1}}\right)$
$\ll\#\{w\mod{c_{2}}\ :\ (w(w+q),c)=1,\ F^{\prime}(w)\equiv 0\mod{c_{1}}\}.$ (4.15)
Since each $\alpha_{j,2}$ is equal to $\alpha_{j,1}$ or $\alpha_{j,1}+1$, any $w$ with $1\leq w\leq c_{2}$ can be written as $w=w_{0}+w_{1}c_{1}$, where $1\leq w_{0}\leq c_{1}$ and $0\leq w_{1}\leq p_{1}\cdots p_{k}$. Moreover, with this notation,
$F^{\prime}(w)\equiv F^{\prime}(w_{0})\pmod{c_{1}}$.
As a result,
$K_{2}(a_{1},a_{2},a_{3},q,c)$
$\qquad\ll\#\{w\mod{c_{1}}\ :\ (w(w+q),c)=1,\ F^{\prime}(w)\equiv 0\mod{c_{1}}\}$ (4.16)
with the implied constant depending on $p_{1},\dots,p_{k}$. This completes the proof of the lemma, since
$F^{\prime}(w)$ $\equiv a_{1}-2a_{2}\overline{w}^{3}-2a_{3}\overline{(w+q)}^{3}\mod{c_{1}}.\qed$
Using Lemma 4.7 in conjunction with a bound on the number of cubic residues (Lemma 3.2), we now give two bounds for $K_{2}$ in the forms required for our application. First, we give a simple pointwise bound needed in the proof of Proposition 2.3.
Lemma 4.8.
Suppose $a,m\in\mathbb{Z}$, $b\geq 2$ and $N\geq 1$ with $(a,b)=1$. Then for any divisor $c$ of $b^{N}$, we have
$K_{2}(m,a,c)=O_{b}(1).$
Proof.
Applying Lemma 4.7 with the notation therein,
$|K_{2}(m,a,c)|$ $\ll_{b}\#\{w\!\!\!\mod{c_{1}}:(w,c)=1,\ m\equiv 2a\overline{w}^{3}\mod{c_{1}}\}.$ (4.17)
If $c_{1}$ is odd, then this reduces further to
$|K_{2}(m,a,c)|\ll_{b}\#\{w\!\!\!\mod{c_{1}}:(w,b)=1,\ m\overline{2a}\equiv\overline{w}^{3}\mod{c_{1}}\}$
and the desired result follows from Lemma 3.2. Now suppose $c_{1}$ is even. If $m$ is odd, then the right-hand side of (4.17) is zero. Then, if $m$ is even, (4.17) can be bounded further as
$|K_{2}(m,a,c)|$ $\ll_{b}\#\{w\!\!\!\mod{c_{1}}:(w,c)=1,\ (m/2)\overline{a}\equiv\overline{w}^{3}\mod{c_{1}/2}\}$
and the desired result again follows from Lemma 3.2. ∎
Next, we give an averaged bound required for Proposition 2.4 which arises after we apply the $B$-process (Lemma 4.1).
Lemma 4.9.
Suppose $a,m\in\mathbb{Z}$, $b\geq 2$ and $N\geq 1$ with $(a,b)=1$. Then for any divisor $c$ of $b^{N}$, we have
$\sum_{|q|\leq Q}|K_{2}(m,a,-a,q,c)|\ll_{b}Q+b^{N/2}.$
Proof.
With the notation of Lemma 4.7,
$\sum_{\begin{subarray}{c}|q|\leq Q\end{subarray}}|K_{2}(m,a,-a,q,c)|\ll_{b}S_{1},$ (4.18)
where $S_{1}$ counts the number of solutions to
$2a(\overline{w}^{3}-\overline{(w+q)}^{3})\equiv m\!\!\mod{c_{1}}$
in variables $w,q$ satisfying
$0\leq w<c_{1},\quad|q|<Q.$
For each fixed $w\!\!\mod{c_{1}}$, we reduce $q\!\!\mod{c_{1}}$ then apply the change of variable
$q\rightarrow q-w.$
This implies that
$S_{1}\ll\left(1+\frac{Q}{c_{1}}\right)S_{2},$ (4.19)
where $S_{2}$ counts the number of solutions to
$2a(w^{3}-q^{3})\equiv m\!\!\mod{c_{1}}$
with
$0\leq w<c_{1},\quad 0\leq q<c_{1},\quad\gcd(wq,c_{1})=1$
and we have made the further change of variables
$w\rightarrow\overline{w},\quad q\rightarrow\overline{q}.$
Now, by Lemma 3.2, we have that
$S_{2}\ll_{b}S_{3}$ (4.20)
where $S_{3}$ counts the number of solutions to
$2a(w-q)\equiv m\!\!\mod{c_{1}}$ (4.21)
in variables
$0\leq w<c_{1},\quad 0\leq q<c_{1},\quad\gcd(wq,c_{1})=1.$
Since we are assuming $(a,b)=1$ and thus $(a,c_{1})=1$, it follows that (4.21) has at most $c_{1}$ solutions, since fixing $w$ uniquely determines $q$. That is,
$S_{2}\ll_{b}S_{3}\ll c_{1}\leq b^{N/2}.$ (4.22)
The result then follows by combining (4.18), (4.19) and (4.22). ∎
5  Proof of Proposition 2.3
Having established the preliminary results in Section 4, we are now able to prove Proposition 2.3. The general idea is to build on the proof of Proposition 2.2, applying Poisson summation (Lemma 4.2) and relevant estimates for oscillatory integrals and exponential sums from Sections 4.2 and 4.3.
Proof of Proposition 2.3.
Let $\varepsilon>0$. As in the proof Proposition 2.2, it suffices to consider palindromes of a fixed length $N$, showing that
$\{n\in\Pi^{*}_{b}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}\ll_{b,\varepsilon}\frac{b^{2N/3}}{D^{2/3-\varepsilon}}$
for $b^{N/4}\leq D\leq b^{2N/5}$. In what follows, we let $K$ be the largest integer such that $b^{K}\leq b^{N/3}/D^{1/3+\varepsilon}$}, so that in particular,
$b^{K}\asymp_{b}\frac{b^{N/3}}{D^{1/3+\varepsilon}}.$ (5.1)
Following the start of the proof of Proposition 2.2, we have
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$
$\leq b^{K}\#\left\{d\sim D,\ s\in\left[\frac{b^{N-1}}{4D^{2}},\frac{b^{N}}{D^{2}}\right]:d^{2}s\equiv a\ \text{(mod}\ b^{K}\text{)},\ (sd,b)=1,\ \left|d^{2}s-\alpha\right|\leq\frac{b^{N}}{b^{K}}\right\}$ (5.2)
for some integer $a$ with $0\leq a<b^{K}$ that does not end in a 0 in base $b$ and $\alpha$ the $N$-digit number beginning with the mirror image of $a$ followed by a suitable number of trailing zeros. Since we are only considering palindromes in $\Pi^{*}_{b}(N)$ we may also assume that $a$ is coprime to $b$.
We now introduce two smoothing functions into our argument, with valid choices plotted in Figure 1 for clarity. First, let $\psi$ be a smooth, piecewise monotonic function compactly supported on $[1/2,5/2]$ with $\psi(x)=1$ for all $x\in[1,2]$. Second, let $\phi$ be a smooth, piecewise monotonic function compactly supported on [2,2][-2,2] with $\phi(x)=1$ for all $x\in[-1,1]$.
Figure 1. Example choices of the functions $\psi(x)$ and $\phi(x)$, plotted using Desmos [18]. Importantly, each function is smooth, compactly supported, piecewise monotonic and bounded.
Using these smoothing functions, we can bound (5) further as
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$
$\qquad\qquad\qquad\leq b^{K}\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\\ (d,b)=1\end{subarray}}\psi\left(\frac{d}{D}\right)\sum_{\begin{subarray}{c}s\in\mathbb{Z}\\ s\equiv a\overline{d}^{2}\ \text{mod}\ b^{K}\end{subarray}}\phi((s-\alpha/d^{2})D^{2}b^{K-N}),$ (5.3)
where $\overline{d}$ denotes a fixed representative of the residue class of the multiplicative inverse of $d$ mod $b^{K}$. Applying Poisson summation (Lemma 4.2),
$\sum_{\begin{subarray}{c}s\in\mathbb{Z}\\ s\equiv a\overline{d}^{2}\ \text{mod}\ b^{K}\end{subarray}}\phi((s-\alpha/d^{2})D^{2}b^{K-N})=$
$\quad\quad\quad\quad\quad\frac{1}{b^{K}}\sum_{\ell\in\operatorname{\mathbb{Z}}}e\left(\frac{a\overline{d}^{2}\ell}{b^{K}}\right)\int_{\operatorname{\mathbb{R}}}\phi((u-\alpha/d^{2})D^{2}b^{K-N})e\left(-\frac{u\ell}{b^{K}}\right)\mathrm{d}u.$
Performing a linear change of variables in the above integral then gives
$\int_{\operatorname{\mathbb{R}}}\phi((u-\alpha/d^{2})D^{2}b^{K-N})e\left(-\frac{u\ell}{b^{K}}\right)\mathrm{d}u=\frac{b^{N}}{D^{2}b^{K}}e\left(-\frac{\alpha\ell}{d^{2}b^{K}}\right)\widehat{\phi}\left(\frac{\ell b^{N}}{D^{2}b^{2K}}\right).$
Substituting back into (5), we have
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$
$\qquad\qquad\ll\frac{b^{N}}{D^{2}b^{K}}\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\\ (d,b)=1\end{subarray}}\psi\left(\frac{d}{D}\right)\sum_{\ell\in\operatorname{\mathbb{Z}}}e\left(-\frac{\alpha\ell}{d^{2}b^{K}}\right)e\left(\frac{a\overline{d}^{2}\ell}{b^{K}}\right)\widehat{\phi}\left(\frac{\ell b^{N}}{D^{2}b^{2K}}\right).$ (5.4)
Consider the Fourier transform $\widehat{\phi}$. Applying integration by parts $k\geq 1$ times, one has for any $x\in\operatorname{\mathbb{R}}/\{0\}$
$\widehat{\phi}(x)=\int_{\operatorname{\mathbb{R}}}\phi(y)e(-xy)\mathrm{d}y=\left(\frac{1}{2\pi i}\right)^{k}\frac{1}{x^{k}}\int_{\operatorname{\mathbb{R}}}\phi^{(k)}(y)e(-xy)\mathrm{d}y\ll\frac{1}{|x|^{k}},$
noting that $\phi^{(k)}(y)$ is bounded on $\mathbb{R}$ since $\phi$ is smooth and compactly supported in [2,2][-2,2]. This implies that
$\sum_{|\ell|>D^{2}b^{2K}/b^{N(1-\varepsilon)}}e\left(-\frac{\alpha\ell}{d^{2}}\right)e\left(\frac{a\overline{d}^{2}\ell}{b^{K}}\right)\widehat{\phi}\left(\frac{\ell b^{N}}{D^{2}b^{2K}}\right)$ $\ll\sum_{|\ell|>D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left|\widehat{\phi}\left(\frac{\ell b^{N}}{D^{2}b^{2K}}\right)\right|$
$\ll\sum_{|\ell|>D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left(\frac{D^{2}b^{2K}}{|\ell|b^{N}}\right)^{k}$
$\ll b^{N\varepsilon(1-k)}=O(b^{-1000N})$
after taking $k$ sufficiently large, depending only on $\varepsilon$. Hence, writing
$\sum_{\ell\in\operatorname{\mathbb{Z}}}=\sum_{\ell=0}+\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}+\sum_{|\ell|>D^{2}b^{2K}/b^{N(1-\varepsilon)}}$
we have that (5) is further bounded as
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}\ll\frac{b^{N}}{Db^{K}}+\frac{b^{N}}{D^{2}b^{K}}\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left|S(\ell)\right|$ (5.5)
where
$S(\ell)=\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\\ (d,b)=1\end{subarray}}\psi\left(\frac{d}{D}\right)e\left(-\frac{\alpha\ell}{d^{2}b^{K}}\right)e\left(\frac{a\overline{d}^{2}\ell}{b^{K}}\right).$ (5.6)
We now partition the sum in (5.5) according to the greatest common divisor of $b^{K}$ and $\ell$:
$\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left|S(\ell)\right|=\sum_{\begin{subarray}{c}c|b^{K}\end{subarray}}\tilde{S}(c)$ (5.7)
where
$\tilde{S}(c)=\sum_{\begin{subarray}{c}1\leq|\ell|\leq D^{2}b^{2K}c/b^{N(1-\varepsilon)+K}\\ (\ell,b)=1\end{subarray}}\left|S(b^{K}\ell/c)\right|.$ (5.8)
The condition $(\ell,b)=1$ is important here, as it will ultimately allow us to apply Lemma 4.8. Now, consider
$S(b^{K}\ell/c)=\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\\ (d,b)=1\end{subarray}}\psi\left(\frac{d}{D}\right)e\left(-\frac{\alpha\ell}{d^{2}c}\right)e\left(\frac{a\overline{d}^{2}\ell}{c}\right).$
By Lemma 4.2
$S(b^{K}\ell/c)$ $=\frac{1}{c^{1/2}}\sum_{m\in\operatorname{\mathbb{Z}}}K_{2}(m,a\ell,c)\int_{\operatorname{\mathbb{R}}}\psi\left(\frac{u}{D}\right)e\left(-\frac{\alpha\ell}{u^{2}c}-\frac{mu}{c}\right)\mathrm{d}u$
$=\frac{D}{c^{1/2}}\sum_{m\in\operatorname{\mathbb{Z}}}K_{2}(m,a\ell,c)\int_{\operatorname{\mathbb{R}}}\psi\left(u\right)e\left(-\frac{\alpha\ell}{D^{2}u^{2}c}-\frac{mDu}{c}\right)\mathrm{d}u.$ (5.9)
We next truncate summation by considering the decay of the integral in (5.9). For fixed $m$, we write
$F(u)=-\frac{\alpha\ell}{u^{2}D^{2}c}-\frac{mDu}{c}$
for the phase of the exponential in (5.9). Here, we restrict to $u\in[1/2,5/2]$ since this is the support of $\psi$. Now,
$F^{\prime}(u)$ $=\frac{2\alpha\ell}{u^{3}D^{2}c}-\frac{mD}{c},$ (5.10)
$|F^{(k)}(u)|$ $=\frac{(k+1)!\thinspace\alpha|\ell|}{u^{k+2}D^{2}c}\qquad k\geq 2.$ (5.11)
If $|m|>\alpha|\ell|/D^{3-\varepsilon}$ then the second term in (5.10) dominates. That is,
$|F^{\prime}(u)|\asymp_{\varepsilon}\frac{|m|D}{c},\qquad\text{for}\ |m|>\frac{\alpha|\ell|}{D^{3-\varepsilon}}.$
Thus, by Lemma 4.3, for any integer $k\geq 2$
$\left|\int_{\operatorname{\mathbb{R}}}\psi(u)e\left(-\frac{\alpha\ell}{u^{2}D^{2}c}-\frac{rDu}{c}\right)\mathrm{d}u\right|\ll_{k,\varepsilon}\left(\frac{c}{|m|D}\right)^{k}\quad\text{if}\quad|m|>\frac{\alpha|\ell|}{D^{3-\varepsilon}}.$
Since $c\leq b^{K}$, we get
$\sum_{|m|>\alpha|\ell|/D^{3-\varepsilon}}\left|\int_{\operatorname{\mathbb{R}}}\psi(u)e\left(-\frac{\alpha\ell}{u^{2}D^{2}c}-\frac{mDu}{c}\right)\mathrm{d}u\right|$ $\ll_{k,\varepsilon}\frac{b^{Kk}}{D^{k}}\sum_{|m|>\alpha|\ell|/D^{3-\varepsilon}}\frac{1}{|m|^{k}}$
$\ll_{k}\frac{b^{Kk}D^{(3-\varepsilon)(k-1)}}{D^{k}\alpha^{k-1}|\ell|^{k-1}}.$ (5.12)
Recalling that $|\ell|\geq 1$, $\alpha\asymp_{b}b^{N}$, $D\leq x^{0.4}$ and $b^{K}\leq b^{N/3}/D^{1/3}$, we have that (5) can be further bounded as
$\sum_{|m|>\alpha|\ell|/D^{3-\varepsilon}}\left|\int_{\operatorname{\mathbb{R}}}\psi(u)e\left(-\frac{\alpha\ell}{u^{2}D^{2}c}-\frac{mDu}{c}\right)\mathrm{d}u\right|\ll_{k,\varepsilon,b}\frac{1}{b^{N(1+\varepsilon(k-1))}}.$ (5.13)
Hence, upon taking $k$ to be sufficiently large, we have
$S(b^{K}\ell/c)$
$=\frac{D}{c^{1/2}}\sum_{|m|\leq\alpha|\ell|/D^{3-\varepsilon}}K_{2}(m,a\ell,c)\int_{\operatorname{\mathbb{R}}}\psi\left(u\right)e\left(\frac{\alpha\ell}{D^{2}u^{2}c}-\frac{mDu}{c}\right)\mathrm{d}u+O_{b,\varepsilon}(b^{-100N}).$ (5.14)
By (5.11) and Lemma 4.6
$\int_{\operatorname{\mathbb{R}}}\psi\left(u\right)e\left(\frac{\alpha\ell}{D^{2}u^{2}c}-\frac{mDu}{c}\right)\mathrm{d}u\ll\frac{Dc^{1/2}}{(\alpha|\ell|)^{1/2}},$
which substituted into (5) implies
$S(b^{K}\ell/c)$ $\ll\frac{D^{2}}{(\alpha|\ell|)^{1/2}}\sum_{|m|\leq\alpha|\ell|/D^{3-\varepsilon}}|K_{2}(m,a\ell,c)|+O_{b,\varepsilon}(b^{-100N}).$
Recalling that $(a\ell,b)=1$, Lemma 4.8 gives $K_{2}(m,a\ell,c)=O_{b}(1)$ and thus
$S(b^{K}\ell/c)\ll_{b,\varepsilon}\frac{(\alpha|\ell|)^{1/2}}{D^{1-\varepsilon}}.$ (5.15)
Substituting (5.15) back into (5.8) and (5.7), and noting that $\alpha\asymp_{b}b^{N}$ then gives
$\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left|S(\ell)\right|\ll_{b,\varepsilon}d(b^{K})\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\frac{(\alpha|\ell|)^{1/2}}{D^{1-\varepsilon}}\ll_{b,\varepsilon}\frac{D^{2+\varepsilon}b^{(3+\varepsilon)K}}{b^{N(1-\frac{3}{2}\varepsilon)}},$ (5.16)
where $d(b^{K})$ is the number of divisors of $b^{K}$, for which one has the standard bound $d(b^{K})\ll_{\varepsilon}b^{\varepsilon K}$. Finally, substituting (5.16) into (5.5) and suitably rescaling $\varepsilon$ yields
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$ $\ll_{b,\varepsilon}\frac{b^{N}}{Db^{K}}+D^{3\varepsilon}b^{2K}.$
Recalling the choice (5.1) of $b^{K}$, we complete the proof. ∎
6  Proof of Proposition 2.4
In this final section, we prove Proposition 2.4. As discussed in Section 2, upon combination with Propositions 2.2 and 2.3 this allows us to deduce our main result, Theorem 1.2.
The proof of Proposition 2.4 begins in the same way as that of Proposition 2.3. However, this time we also apply our $A$-process (Lemma 4.1), which requires different exponential sum estimates and notably allows us to attain a non-trivial bound on $S_{b}(x,D)$ for values of $D$ less than $x^{1/4}$.
Proof of Proposition 2.4.
Let $\varepsilon>0$. As in the proofs of Propositions 2.2 and 2.3, it suffices to consider palindromes of a fixed length $N$, showing that
$\{n\in\Pi^{*}_{b}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}\ll_{b,\varepsilon}\frac{b^{7N/11}}{D^{13/22-\varepsilon}}.$
for $b^{3N/13}\leq D\leq b^{8N/31-\varepsilon}$. In what follows, we let $K$ be the largest integer such that $b^{K}\leq b^{4N/11}/D^{9/22+\varepsilon}$}, so that in particular,
$b^{K}\asymp_{b}\frac{b^{4N/11}}{D^{9/22+\varepsilon}}.$ (6.1)
We now argue as in the proof of Proposition 2.3 with the same notation up until (5.5) and without the smoothing factor $\psi(d/D)$. This gives
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}\ll\frac{b^{N}}{Db^{K}}+\frac{b^{N}}{D^{2}b^{K}}\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left|S(\ell)\right|,$ (6.2)
where
$S(\ell)=\sum_{\begin{subarray}{c}d\sim D\\ (d,b)=1\end{subarray}}e\left(\frac{\alpha\ell}{d^{2}b^{K}}\right)e\left(-\frac{a\overline{d}^{2}\ell}{b^{K}}\right).$ (6.3)
As in the proof of Proposition 2.3, we partition the sum in (6.2) according to the greatest common divisor of $\ell$ and $b^{K}$:
$\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left|S(\ell)\right|=\sum_{\begin{subarray}{c}c|b^{K}\end{subarray}}\tilde{S}(c)$ (6.4)
where
$\tilde{S}(c)=\sum_{\begin{subarray}{c}1\leq|\ell|\leq D^{2}b^{2K}c/b^{N(1-\varepsilon)+K}\\ \gcd(\ell,b)=1\end{subarray}}\left|S(b^{K}\ell/c)\right|.$ (6.5)
Consider
$S(b^{K}\ell/c)=\sum_{\begin{subarray}{c}d\sim D\\ (d,b)=1\end{subarray}}e\left(-\frac{\alpha\ell}{d^{2}c}\right)e\left(\frac{a\overline{d}^{2}\ell}{c}\right).$
We now diverge from the proof of Proposition 2.3, and apply Lemma 4.1 with
$Q=D^{1/2}.$ (6.6)
This gives
$S(b^{K}\ell/c)\ll D^{3/4}+D^{1/4}\left(\sum_{q\leq D^{1/2}}|\tilde{S}(\ell,q)|\right)^{1/2}$ (6.7)
where
$\tilde{S}(\ell,q)=\sum_{\begin{subarray}{c}d\in\operatorname{\mathbb{Z}}\end{subarray}}\psi\left(\frac{d}{D}\right)e\left(-\frac{\alpha\ell}{c}\left(\frac{1}{d^{2}}-\frac{1}{(d+q)^{2}}\right)\right)e\left(\frac{a\ell(\overline{d}^{2}-\overline{(d+q)}^{2})}{c}\right)$
with $\psi(d/D)$ the same smoothing function from the proof of Proposition 2.3 (see Figure 1). We now apply Lemma 4.2 (Poisson summation) to get
$\tilde{S}(\ell,q)$
$=\frac{1}{c^{1/2}}\sum_{m\in\operatorname{\mathbb{Z}}}K_{2}(m,a\ell,-a\ell,q,c)\int_{\operatorname{\mathbb{R}}}\psi\left(\frac{u}{D}\right)e\left(-\frac{\alpha\ell}{c}\left(\frac{1}{u^{2}}-\frac{1}{(u+q)^{2}}\right)-\frac{mu}{c}\right)\mathrm{d}u$
$=\frac{D}{c^{1/2}}\sum_{m\in\operatorname{\mathbb{Z}}}K_{2}(m,a\ell,-a\ell,q,c)\int_{\operatorname{\mathbb{R}}}\psi\left(u\right)e\left(-\frac{\alpha\ell}{c}\left(\frac{1}{D^{2}u^{2}}-\frac{1}{(Du+q)^{2}}\right)-\frac{mDu}{c}\right)\mathrm{d}u.$ (6.8)
With a view to apply Lemma 4.3, we let
$F(u)=-\frac{\alpha\ell}{c}\left(\frac{1}{D^{2}u^{2}}-\frac{1}{(Du+q)^{2}}\right)-\frac{mDu}{c}$
denote the phase of the exponential in (6.8), with $u\in[1/2,5/2]$ as this is the support of $\psi$. We have
$F^{\prime}(u)$ $=\frac{2\alpha\ell}{c}\left(\frac{1}{D^{2}u^{3}}-\frac{D}{(Du+q)^{3}}\right)-\frac{mD}{c}$
$=\frac{2\alpha\ell}{c}\left(\frac{3D^{2}u^{2}q+3Duq^{2}+q^{3}}{D^{2}u^{3}(Du+q)^{3}}\right)-\frac{mD}{c},$ (6.9)
$|F^{(k)}(u)|$ $=(k+1)!\frac{\alpha|\ell|}{c}\left(\frac{1}{D^{2}u^{k+2}}-\frac{D^{k}}{(Du+q)^{k+2}}\right),\qquad k\geq 2.$ (6.10)
Since $q\leq D^{1/2}$, we see that the second term in (6.9) dominates for $|m|>\alpha|\ell|q/D^{4-\varepsilon}$. Then, since $\alpha\asymp_{b}b^{N}$ and
$1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}$
this implies that
$|F^{\prime}(u)|\asymp_{\varepsilon}\frac{|m|D}{c},\qquad\text{for}\quad|m|>\frac{b^{2K}q}{D^{2-\varepsilon}}b^{N\varepsilon}.$ (6.11)
As we are only considering $D$ with $\log D\asymp_{b}N$, we may suitably rescale $\varepsilon$ to recast and simplify (6.11) as
$|F^{\prime}(u)|\asymp_{b,\varepsilon}\frac{|m|D}{c},\qquad\text{for}\quad|m|>\frac{b^{2K}q}{D^{2-\varepsilon}}.$
Thus, applying Lemma 4.3, we have for any $k\geq 2$,
$\sum_{|m|>b^{2K}q/D^{2-\varepsilon}}\left|\int_{\operatorname{\mathbb{R}}}\psi\left(u\right)e\left(-\frac{\alpha\ell}{c}\left(\frac{1}{D^{2}u^{2}}-\frac{1}{(Du+q)^{2}}\right)-\frac{mDu}{c}\right)\mathrm{d}u\right|$
$\qquad\ll_{k,\varepsilon}\frac{b^{Kk}}{D^{k}}\sum_{|m|>b^{2K}q/D^{2-\varepsilon}}\frac{1}{|m|^{k}}\ll_{k}\frac{qb^{2K}}{D^{2-\varepsilon}}\left(\frac{D^{1-\varepsilon}}{b^{K}q}\right)^{k}\leq\frac{b^{2K}}{D^{3/2-\varepsilon}}\left(\frac{D^{1-\varepsilon}}{b^{K}}\right)^{k},$ (6.12)
where in the last step we used that $1\leq q\leq D^{1/2}$. By our choice (6.1) of $b^{K}$, and the fact that $b^{3N/13}\leq D\leq b^{8N/31-\varepsilon}$, we have
$\frac{D^{1-\varepsilon}}{b^{K}}\leq D^{-2\varepsilon}\leq b^{-\frac{6N}{13}\varepsilon}.$
Hence, taking $k$ to be sufficiently large in terms of $\varepsilon$, we obtain
$\tilde{S}(\ell,q)=\frac{D}{c^{1/2}}\sum_{|m|\leq\frac{b^{2K}q}{D^{2-\varepsilon}}}K_{2}(m,a\ell,-a\ell,q,c)$
$\qquad\qquad\qquad\qquad\cdot\int_{\operatorname{\mathbb{R}}}\psi\left(u\right)e\left(-\frac{\alpha\ell}{c}\left(\frac{1}{D^{2}u^{2}}-\frac{1}{(Du+q)^{2}}\right)-\frac{mDu}{c}\right)\mathrm{d}u$ (6.13)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad+O_{b,\varepsilon}(b^{-100N}).$
Now, from (6.10),
$F^{\prime\prime}(u)\gg\frac{\alpha|\ell|q}{cD^{3}},\quad u\in\left[\frac{1}{2},\frac{5}{2}\right].$
Thus, applying Lemma 4.6,
$\int_{\operatorname{\mathbb{R}}}\psi\left(u\right)e\left(\frac{\alpha\ell}{c}\left(\frac{1}{D^{2}u^{2}}-\frac{1}{(D^{2}u+q)^{2}}\right)-\frac{mDu}{c}\right)\mathrm{d}x\ll\left(\frac{\alpha|\ell|q}{cD^{3}}\right)^{-1/2}$
which substituted into (6) implies that
$\tilde{S}(\ell,q)\ll_{b,\varepsilon}\frac{D^{5/2}}{(\alpha|\ell|q)^{1/2}}\sum_{|m|\leq\frac{b^{2K}q}{D^{2-\varepsilon}}}|K_{2}(m,a\ell,-a\ell,q,c)|.$
Therefore, from (6.7),
$S(b^{K}\ell/c)$ $\ll_{b,\varepsilon}D^{3/4}+\frac{D^{3/2}}{(\alpha|\ell|)^{1/4}}\left(\sum_{q\leq D^{1/2}}\frac{1}{q^{1/2}}\sum_{|m|\leq\frac{b^{2K}q}{D^{2-\varepsilon}}}|K_{2}(m,a\ell,-a\ell,q,c)|\right)^{1/2}$
$\ll D^{3/4}+\frac{D^{3/2}}{(\alpha|\ell|)^{1/4}}\left(\sup_{m\in\mathbb{Z}}\sum_{q\leq D^{1/2}}\frac{b^{2K}q^{1/2}}{D^{2-\varepsilon}}|K_{2}(m,a\ell,-a\ell,q,c)|\right)^{1/2}$
$\ll D^{3/4}+\frac{b^{K}D^{5/8+\varepsilon/2}}{(\alpha|\ell|)^{1/4}}\left(\sup_{m\in\mathbb{Z}}\sum_{q\leq D^{1/2}}|K_{2}(m,a\ell,-a\ell,q,c)|\right)^{1/2}.$
Since $\gcd(a\ell,b)=1$, an application of Lemma 4.9 gives
$\sup_{m\in\operatorname{\mathbb{Z}}}\sum_{q\leq D^{1/2}}|K_{2}(m,a\ell,-a\ell,q,c)|\ll_{b}D^{1/2}+b^{K/2}\ll b^{K/2},$
where we have used that $D^{1/2}\leq b^{K/2}$, which follows from the definition (6.1) of $b^{K}$ and the fact that $D\leq b^{8N/31-\varepsilon}$. Therefore,
$S(b^{K}\ell/c)$ $\ll_{b,\varepsilon}D^{3/4}+\frac{b^{5K/4}D^{5/8+\varepsilon/2}}{(\alpha|\ell|)^{1/4}}.$ (6.14)
Substituting (6.14) back into (6.4) and (6.5) then yields
$\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left|S(\ell)\right|$ $\ll_{b,\varepsilon}d(b^{K})\sum_{1\leq|\ell|\leq D^{2}b^{2K}/b^{N(1-\varepsilon)}}\left(D^{3/4}+\frac{b^{5K/4}D^{5/8+\varepsilon/2}}{(\alpha|\ell|)^{1/4}}\right)$
$\ll_{b,\varepsilon}b^{\varepsilon K}\left(\frac{D^{11/4}b^{2K}}{b^{N(1-\varepsilon)}}+\frac{b^{11K/4}D^{17/8+\varepsilon/2}}{b^{N(1-3\varepsilon/4)}}\right),$
where we have used that $\alpha\asymp_{b}b^{N}$ and $d(b^{K})\ll_{\varepsilon}b^{\varepsilon K}$. From here, (6.2) gives that
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}\ll_{b,\varepsilon}\frac{b^{N}}{Db^{K}}+b^{K}D^{3/4+\varepsilon}+b^{7K/4}D^{1/8+\varepsilon},$ (6.15)
where we have suitably rescaled $\varepsilon$. Our choice of $b^{K}$ in (6.1) essentially balances the first and third terms in (6.15), giving
$\#\{n\in\Pi_{b}^{*}(N):\exists d\sim D\ \text{s.t.}\ d^{2}\mid n\}$ $\ll_{b,\varepsilon}\frac{b^{7N/11}}{D^{13/22+\varepsilon}}+b^{4N/11}D^{15/44}.$ (6.16)
Since $b^{3N/13}\leq D\leq b^{8N/31-\varepsilon}$, the second term in (6.16) is of smaller order than the first, and can thus be removed. This completes the proof of the Proposition. ∎